A company has determined that the price and the monthly demand of one of its products are related by the equation

ZOJFumK2GAcAAAAASUVORK5CYII=

where p is the price per unit in dollars and D is the monthly demand. The associated fixed costs are $1,125/month, and the variable costs are $100/unit. Use this information to answer Problem 2-52 and Problem 2-53. Select the closest answer.

2-52 What is the optimal number of units that should be produced and sold each month?

  1. 10 units
  2. 15 units
  3. 20 units
  4. 25 units

2-53 Which of the following values of D represents the breakeven point?

  1. 10 units
  2. 15 units
  3. 20 units
  4. 25 units
Answer #1

92 D= Lyoorp 0² = 4oorp - P=40o-D²= 4oord? TREPO =400o-03 MRE 400-307 Rofit is mac Mrame 400-302=100 - 300=302 0²=100 id=607